OpenStudy (anonymous):

Find the volume and total surface area of a regular tetrahedron whose measure of one edge is 10cm

OpenStudy (anonymous):

do you know the formulaes for the surface area of a tetrahedron?

OpenStudy (anonymous):

I assume not so here it is: \[\sqrt{3} a ^{2}\]

OpenStudy (anonymous):

@bozybosman then you multiply that to 4 right to get the total surface area?

OpenStudy (anonymous):

cry

OpenStudy (anonymous):

there is different ways you can find the surface area and the volume 1. find the base plane are which is [(Sq root 3 ) /4 ] a^2 a= to the edge length 2. find the surface area which is 4( base plane area) or (Sq root 3) a^2 3. To find the volume you can use the formula V= [(Sq root 2) / 12] a^3 I hope this helps

OpenStudy (anonymous):

no so that is the formula for surface area you can derive it from the surface area for a right angled triangle. the formula for the surface area of a single face is: \[\frac{ \sqrt{3} }{ 4 }a ^{2}\]

OpenStudy (anonymous):

1) V = 1/12 (a^3) x 3^(1/2) = 1/12 ( 10^3) x 3^(1/2) = 1000/12 x 3^(1/2) = 144.338 cm^3 (3 dec.plcs)

OpenStudy (anonymous):

how about the Total surface area?

OpenStudy (anonymous):

Total surface area is:\[Total Surface Area =\frac{\sqrt{3}}{4}a^2\]

OpenStudy (anonymous):

a is the apothem? or the edge?

OpenStudy (anonymous):

a is the edge because this is a regular tetrahedron it is the same for all sides